In whole numbers, x + 8 = 12 – 4 has                            solution.
In whole numbers, x + 8 = 12 – 4 has                            solution.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given integers, x + 8 = 12–4.

x + 8 =8
Transpose 8 to RHS then it becomes - 8.

x= 8 - 8

x= 0
In natural numbers, x + 8 = 12–4 has one solution.

Thanks